Please help me solve this:
Find the equation of the line that passes through (1,3) and parallel to 2x-3y=5

Answer by ewatrrr(24785)

 
Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
Find the equation of the line that passes through (1,3) and parallel to 2x-3y=5
2x-3y=5 OR y = (2/3)x - 5/3 m = 2/3
Parallel Lines have equal slopes:
New Line
y = (2/3)x + b ||Using Pt(1,3) to solve for b
3 = 2/3 + b
7/3 = b
y = (2/3)x + 7/3 or 2x-3y = -7